A grocer prepares a mixture of two pastas to sell. For the mixture he uses penne pasta that sells for $4.75 per pound and rigatoni that sells for $6.00 per pound. If he uses 2 pounds of penne pasta and 3 pounds of rigatoni for a 5 pounds mix, how much should he sell the pasta for per pound?
Answer by checkley77(12844)

3*4.75+2*6=5x
14.25+12=5x
5x=26.25
x=26.25/5
x=$5.25 per pound for the mix.